OpenStudy (anonymous):

how to solve x^2+kx-19

OpenStudy (anonymous):

it's not an equation, it can't be solve... maybe a "=0" is missing ?

OpenStudy (anonymous):

im suposee to find the value of k

OpenStudy (anonymous):

yeah but you should have an "= something"

OpenStudy (amistre64):

since c=sqrt(k/2) id start with that :)

OpenStudy (amistre64):

i assume it wants you to solve for k such that it is a complete square?

OpenStudy (amistre64):

c=(k/2)^2 that is

OpenStudy (amistre64):

had a synapse crossed lol

OpenStudy (anonymous):

i don't understand... what is c here ?

OpenStudy (amistre64):

ax^2 +bx + c is the usual generic format for a quad

OpenStudy (amistre64):

a complete square is then the case where c = (b/2)^2

OpenStudy (anonymous):

yeah, but c = 19... ?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so you decide it's a complete square, and then you resolve the thing ?

OpenStudy (amistre64):

thats my assumption; i could be wrong since the information in the problem is open to interpretation :)

OpenStudy (anonymous):

okay, i understand know. i absolutely didn't know that thing of "complete square". must not being teached by french teachers.

OpenStudy (amistre64):

.... well, you cant blame them to much .. they are french :)

OpenStudy (anonymous):

hey, i'm french too :p